OpenStudy (anonymous):

Two trains leave the station at the same time, one heading east and the other west. The eastbound train travels 12 miles per hour faster than the westbound train. If the two trains are 800 miles apart after 4 hours, what is the rate of the eastbound train? Do not do any rounding.

OpenStudy (anonymous):



OpenStudy (anonymous):

Just use relative Velocity theory here..))

OpenStudy (anonymous):

12 + x + x = 12 + 2x === Speed Distance ==== 800 T = 4 S = D/T 12 + 2x = 800/4 12 + 2x = 200 2x = 188 x = 94 miles/h